# PRO-V-R1: Reasoning Enhanced Programming Agent for RTL Verification

### v3

@@ 2.1. LLM-based Verilog generation and verification
$2.$ Tool-Enhanced LLM Agents with Post-training
@@ 3.1. Agentic Verification System
Testbench Generation and Functional Modeling. The FRM Generator (Fig. 3 (b)) takes the JSON Input Signal from step  1 and the RTL SPEC to construct a testbench.json artifact, pairing input vectors with cycle-aligned expected outputs. To achieve this, the agent leverages LLM-based code generation (Sec. 2.2) to sample NN Python-based FRMs acting as “golden” behavioral oracles. These scripts are executed by the Code Interpreter to compute NN sets of expected outputs for the step  1 testcases. This program-aided approach circumvents LLM limitations in direct arithmetic (gao2023pal; li2024gsmplus) by applying precise, executable logic.
@@
Iterative Refinement and Verification. The Judge agent (Fig. 3 (b)) receives the NN candidate artifacts from step  2 and drives a self-improvement loop (Fig. 3 (c)) combining majority voting with verification. First, the system selects the most consistent candidate via majority voting. The Judge then validates the selected outputs; if errors are detected, it performs targeted corrections by invoking the Testbench Edit Tool. As detailed in Tab. 1, this tool allows the agent to programmatically add, replace, or remove entries in testbench.json (e.g., rectifying a single failing vector). Upon validation or budget exhaustion, the final testbench.json is passed to the Simulation Driver to compile and orchestrate the Verilator simulation. This driver validates the DUT by comparing its actual execution results—driven by the stimuli from step  1 —against the expected reference outputs encapsulated in the testbench.json.
@@ 3.2. Agentic training framework for RTL verification
This filter ensures the agentic flow successfully generates a runnable test harness (composed of $C_{TC}$ and $C_{FRM}$) and that the golden RTL’s simulation output precisely matches the FRM’s output for all testcases $C_{TC}$. This process validates the functional correctness of the generated $C_{FRM}$. We then collect these verified interactions to form our curated SFT dataset, $\mathcal{D}_{SFT}$. Each data point $d\in\mathcal{D}_{SFT}$ is a trajectory composed of components for each agent, structured as: $(P_{i},T_{i},C_{i}),\text{where}\,i\in\{\text{TC},\text{FRM}\}$. Here, $P_{i}$ represents the prompt, $T_{i}$ is the reasoning trace, and $C_{i}$ is the generated code for agent ii.
@@
  • TC Agent Functional Reward ($R_{func}^{\text{TC}}$): This is a binary reward for generating a runnable harness. It is 11 if $C_{TC}$ and $C_{FRM}$ are syntactically valid and can be successfully compiled and simulated together, and 0 otherwise.
  • FRM Agent Functional Reward ($R_{func}^{\text{FRM}}$): This is the primary verification success reward. It is 11 only if all test cases pass verification—that is, the strict filter $\mathcal{V}(\cdot)$ (defined in Eq. (3.2)) evaluates to true. This reward is contingent on the harness being runnable (i.e., $R_{func}^{\text{TC}}=1$).

### v4

@@ 2.1. LLM-based Verilog generation and verification
2.2. Tool-Enhanced LLM Agents with Post-training
@@ 3.1. Agentic Verification System
Testbench Generation and Functional Modeling. The FRM Generator (Fig. 3 (b)) takes the JSON Input Signal from step  1 and the RTL SPEC to construct a testbench.json artifact, pairing input vectors with cycle-aligned expected outputs. To achieve this, the agent leverages LLM-based code generation (Sec. 2.2) to sample $N$ Python-based FRMs acting as “golden” behavioral oracles. These scripts are executed by the Code Interpreter to compute $N$ sets of expected outputs for the step  1 testcases. This program-aided approach circumvents LLM limitations in direct arithmetic (gao2023pal; li2024gsmplus) by applying precise, executable logic.
@@
Iterative Refinement and Verification. The Judge agent (Fig. 3 (b)) receives the $N$ candidate artifacts from step  2 and drives a self-improvement loop (Fig. 3 (c)) combining majority voting with verification. First, the system selects the most consistent candidate via majority voting. The Judge then validates the selected outputs; if errors are detected, it performs targeted corrections by invoking the Testbench Edit Tool. As detailed in Tab. 1, this tool allows the agent to programmatically add, replace, or remove entries in testbench.json (e.g., rectifying a single failing vector). Upon validation or budget exhaustion, the final testbench.json is passed to the Simulation Driver to compile and orchestrate the Verilator simulation. This driver validates the DUT by comparing its actual execution results—driven by the stimuli from step  1 —against the expected reference outputs encapsulated in the testbench.json.
@@ 3.2. Agentic training framework for RTL verification
This filter ensures the agentic flow successfully generates a runnable test harness (composed of $C_{TC}$ and $C_{FRM}$) and that the golden RTL’s simulation output precisely matches the FRM’s output for all testcases $C_{TC}$. This process validates the functional correctness of the generated $C_{FRM}$. We then collect these verified interactions to form our curated SFT dataset, $\mathcal{D}_{SFT}$. Each data point $d\in\mathcal{D}_{SFT}$ is a trajectory composed of components for each agent, structured as: $(P_{i},T_{i},C_{i}),\text{where}\,i\in\{\text{TC},\text{FRM}\}$. Here, $P_{i}$ represents the prompt, $T_{i}$ is the reasoning trace, and $C_{i}$ is the generated code for agent $i$.
@@
  • TC Agent Functional Reward ($R_{func}^{\text{TC}}$): This is a binary reward for generating a runnable harness. It is $1$ if $C_{TC}$ and $C_{FRM}$ are syntactically valid and can be successfully compiled and simulated together, and 0 otherwise.
  • FRM Agent Functional Reward ($R_{func}^{\text{FRM}}$): This is the primary verification success reward. It is $1$ only if all test cases pass verification—that is, the strict filter $\mathcal{V}(\cdot)$ (defined in Eq. (3.2)) evaluates to true. This reward is contingent on the harness being runnable (i.e., $R_{func}^{\text{TC}}=1$).
@@ 5. Conclusion
Acknowledgements.
We thank NVIDIA for providing access to and support for cloud services to perform model training.